5-7x>3
(1-t)/2<(3t-7)/3
/3-2x/=7

Answer by stanbon(57361)

5-7x>3
Rewrite as:
7x<5-3
7x<2
x<2/7
-----------
(1-t)/2<(3t-7)/3
Think something is missing in this post.
-------------
/3-2x/=7
3-2x must be 7 or it must be -7
If 3-2x=7 then 2x=3-7 then x=-2
If 3-2x=-7 then 2x=3+7 then x=5
Cheers,
Stan H